OpenStudy (anonymous):

Evaluate : 3!/2 - 4!/3 + 5!/4 - 6!/5 + .... + 2013!/2014 - 2014!/2013

OpenStudy (empty):

Shouldn't the second to last term really be 2013!/2012 ?

OpenStudy (anonymous):

Oopsss..sorry i was typo there. Yeah the second last should be 2013!/2012

OpenStudy (empty):

Fun problem I'll say that! I don't know where to begin hmm...

OpenStudy (loser66):

Another approaching \(a_1= \dfrac{3!}{2}\\a_3=\dfrac{5!}{4}\\a_5=\dfrac{7!}{6}\\--------\) while \(a_2= -\dfrac{4!}{3}\\a_4=-\dfrac{6!}{5}\\a_6=-\dfrac{8!}{7}---------\)

OpenStudy (loser66):

So our sequence is \(a_n= (-1)^{n+1}\dfrac{(n+2)!}{n+1}\)

OpenStudy (loser66):

I give up!! hehehe.. it's above my head!!

OpenStudy (empty):

What you deleted everything AND gave up?!

OpenStudy (empty):

Damn it

OpenStudy (anonymous):

hey! what do you need help with?

OpenStudy (anonymous):

need a way and solution :)

OpenStudy (anonymous):

Isn't that \[(-1)^{n+1} (n!+(n+1)!)\]

OpenStudy (anonymous):

Telescopes quite nicely

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

So what's next @mukushla ?

ganeshie8 (ganeshie8):

do you see see how \(\dfrac{(n+2)!}{n+1} \) simplifies to \(n! + (n+1)!\) ?

OpenStudy (anonymous):

(n+1)!/n = (n+1)n(n-1)!/n = (n+1)(n-1)! hmmm...

ganeshie8 (ganeshie8):

\[\begin{align}\dfrac{(n+2)!}{n+1} &= \dfrac{(n+2)(n+1)n!}{n+1} \\~\\ &= (n+2)n! = (\color{blue}{n+1}+1)n!\\~\\ & = (n+1)n! + n! \\~\\ &= (n+1)!+n!\end{align}\]

OpenStudy (anonymous):

ah yes i see now :) sorry i look n+1 but should be n+2. Whats next ?

ganeshie8 (ganeshie8):

familiar with sigma notation ? \(\sum\)

ganeshie8 (ganeshie8):

the given sum is same as : \[ [(1+1)!+1!] - [(2+1)!+2!] + [(3+1)!+3!] -\cdots -[(2012+1)!+2012!] \]

OpenStudy (loser66):

Wooooooooooooooah!! it is nice.

ganeshie8 (ganeshie8):

or \[ [2!+1!] - [3!+2!] + [4!+3!] -\cdots -[2013!+2012!] \]

ganeshie8 (ganeshie8):

please medal loser/mukushla, not meh

OpenStudy (loser66):

You work, why medal me?

OpenStudy (anonymous):

Looks the series is nice but not sure which numbers can be cancels ?

ganeshie8 (ganeshie8):

i didnt use my brain, i just used ur work for general term and mukushla's idea of telescoping

OpenStudy (loser66):